Exercise 19-13 EPS; treasury stock; new shares; stock dividends; two years [LO19-5, 19-6]

The Alford Group had 190,000 shares of common stock outstanding at January 1, 2018. The following activities affected common shares during the year. There are no potential common shares outstanding.
  

2018
Feb. 28 Purchased 15,000 shares of treasury stock.
Oct. 31 Sold the treasury shares purchased on February 28.
Nov. 30 Issued 60,000 new shares.
Dec. 31 Net income for 2018 is $1,665,000.
2019
Jan. 15 Declared and issued a 2-for-1 stock split.
Dec. 31 Net income for 2019 is $1,665,000.


Required:

1. Determine the 2018 EPS.
2. Determine the 2019 EPS.
3. At what amount will the 2018 EPS be presented in the 2019 comparative financial statements?

(For all requirements, Enter your answers in thousands.)
Answer #1

(1) 2017:

Numerator: Net Income = $1665000; Preferred dividends = $0

Denominator: Weighted average no of shares

1/1 – 12/31 190000 x (12/12) = 190000

2/28 – 12/31 (15000)x (10/12) = (12500)

10/31 – 12/31 15000 x (2/12) = 2500

11/30 – 12/31 60000 x (1/12) = 5000

Weighted average # shares outstanding 185000

2017 EPS = $1665000 ÷ 185000 = $9

(2) 2018:

Numerator: Net Income = $1665000; Preferred dividends = $0

Weighted average # of shares

1/1 – 12/31 190000–15000+15000+60000 = 250000 x 2 = 500000

2018 EPS = $1665000 ÷ 500000 = $3.33

(3) For comparative purposes, the 2-for-1 stock split in 2018 needs to be retroactively applied to the earliest date being presented, i.e., January 1, 2017.

Weighted Average no of shares outstanding from 2017 = 185000

Retroactive application of stock split x 2

Weighted Average # shares for 2017 = (185000 * 2) = 370000

For comparative purposes:

2017 EPS = $1665000 ÷ 370000 = $4.5

2018 EPS = $1665000 ÷ 500000 = $3.33